# Hire Your Anthropologist! Rethinking Culture Benchmarks Through an Anthropological Lens

### v1

@@ Abstract
Cultural evaluation of large language models has become increasingly important, yet current benchmarks often reduce culture to static facts or homogeneous values. This view conflicts with anthropological accounts that emphasize culture as dynamic, historically situated, and enacted in practice. To analyze this gap, we introduce a four-part framework that categorizes how benchmarks frame culture, such as knowledge, preference, performance, or bias. Using this lens, we qualitatively examine 20 cultural benchmarks and identify six recurring methodological issues, including treating countries as cultures, overlooking within-culture diversity, and relying on oversimplified survey formats. Drawing on established anthropological methods, we propose concrete improvements: incorporating real-world narratives and scenarios, involving cultural communities in design and validation, and evaluating models in context rather than isolation. Our aim is to guide the development of cultural benchmarks that go beyond static recall tasks and more accurately capture the responses of the models to complex cultural situations.
@@
Mai AlKhamissi∗,1   Yunze Xiao∗,1   Badr AlKhamissi2   Mona Diab1 1Carnegie Mellon University  2EPFL {malkhami,yunzex,mdiab}@andrew.cmu.edubadr.alkhamissi@epfl.ch
@@ 1 Introduction
Large language models (LLMs) are now being implemented in translation systems Zhu et al. (2024), educational tools Sonkar et al. (2024), search engines Ziems et al. (2023), and generative platforms that engage diverse public opinions across cultural, linguistic, and political contexts Ziems et al. (2023). As these models become more central to knowledge production and everyday decision-making, it is crucial to assess their sensitivity to cultural nuance. In response, NLP researchers are constructing benchmarks that aim to measure cultural competence: the ability of models to respond appropriately to regionally specific norms, moral frameworks, idioms and socio-political identities.
@@
Figure 1: Framing of Culture in NLP. Our taxonomy of cultural framings used in NLP evaluation. Each quadrant reflects a distinct theoretical lens on culture: what it is, how it is expressed, and what it means to measure. This visual framework guides our critique and re-imagination of cultural benchmarks.
@@
However, in this growing body of work, the concept of culture is often treated as a background variable rather than as a central analytical concern. Benchmark designers rarely engage with anthropology, a discipline that has developed rich tools to understand how culture is lived, contested, and transmitted. Anthropologists study culture not as a fixed set of values or national traits but as a dynamic and situated process shaped by history, power, and everyday meaning-making Geertz (1973). This paper offers anthropological insights to analyze 20 cultural benchmarks used in NLP.
The choice to analyze 20 benchmarks reflects a methodological commitment to qualitative, humanistic traditions, as opposed to conducting a large-scale survey over many more benchmarks. This approach foregrounds epistemologies, assumptions, and cultural politics often obscured in large-scale corpus analyses. Rather than prioritizing breadth, it emphasizes interpretive depth, reflexivity, and contextual understanding, aligning with scholarship that highlights the limits of scalable methods for capturing nuance, bias, and power dynamics in benchmarks. Scholars such as Blodgett et al. and voices in ACM interactions (Liao et al., 2024; Widder and Kneese, 2025) stress that meaningful cultural analysis demands slow, interdisciplinary, and critically reflexive engagement, a position this study adopts through its carefully curated sample.
@@
We argue that while these benchmarks address important questions, such as moral diversity, stereotype reproduction, and regional variation—they often rely on reductive or decontextualized models of culture. Consequently, our contribution is as follows: (1) First, we examine how culture is conceptualized within benchmarks, developing a taxonomy of cultural framings used by evaluators. (2) Second, we show how six common design choices oversimplify culture, reducing it to surface traits or stereotypes, and restricting which forms of cultural knowledge are recognized or ignored. (3) Third, we offer recommendations for building benchmarks that reflect culture as a lived experience.
@@ 2.1 What is Culture
Anthropology treats culture as dynamic, contested, and situated in daily practice, not as a static list of traits. Cultural relativism argued that there are no universal hierarchies and that cultural phenomena must be read in their historical and local contexts (Boas, 1911). The interpretive tradition reframed culture as webs of significance that require a thick description of symbols, rituals, and ordinary acts to understand how meaning is produced (Geertz, 1973). Reflexive work also reminds us that culture accounts are authored and positioned, which places the responsibility on researchers to make assumptions explicit (Clifford and Marcus, 1986). Interactional perspectives emphasize that competence is acted upon in specific scenes where roles, audiences, and settings shape what counts as appropriate behavior and language (Goffman, 1959). Taken together, these views highlight plurality and change over time, and they foreground disagreement as an ordinary feature of cultural life rather than noise to be averaged away. Historical notes and pointers to the broader anthropological literature appear in Appendix A.
@@ 2.2 Culture Benchmarks in NLP
NLP research has developed benchmarks to evaluate the handling of cultural-specific knowledge by language models, but these tools often define culture inconsistently and simplistically compared to anthropological scholarship. Although NLP researchers increasingly recognize that linguistic and social norms vary across cultures and are underrepresented in current resources, significant gaps remain between computational approaches and social science perspectives on culture. The purpose of this work is to foster interdisciplinary dialogue between these fields.
@@ 2.3 Critiques within NLP
Core Limitations of Cultural Benchmarks.
Recent work in NLP has exposed fundamental weaknesses in the conceptualization of culture using cultural benchmarks. In the literature, there is a consensus that current evaluation frameworks reduce culture to static survey-derived constructs that privilege Western-centric assumptions while failing to capture dynamic, situated and contested cultural meanings Liu (2024). Hershcovich et al. (2022), identify unequal cultural representation and the tendency for cross-lingual NLP to remain culturally agnostic, advocating for targeted sampling and Distributionally Robust Optimization to prevent cultural generalization from being disproportionately derived from dominant groups. Zhou et al. Zhou et al. (2025) extends these concerns by critiquing the overreliance on nationality-based categorizations and rigid survey methods that flatten dynamic and intersectional cultural identities, proposing more context-sensitive and stereotype-aware datasets to reflect lived cultural realities.
@@
Researchers have attempted to address these limitations through technical interventions, most notably cultural prompting and semantic augmentation. For example, AlKhamissi et al. (2024) proposed anthropological urging to make LLMs better reflect the opinions of digitally underrepresented groups. Similarly, Tao et al. (2024) shows that explicitly instructing LLMs to adopt specific cultural perspectives can significantly reduce Western bias in most countries and territories, although this approach sometimes introduces new biases and rests on the problematic assumption that culture can be cleaned up by demographic targeting. Although these innovations show promise in controlled settings, they fundamentally maintain the flawed premise that culture can be distilled through survey data or prompt engineering, an assumption increasingly questioned within the research community.
Cultural Framework Approaches and Remaining Gaps.
@@
The framework presented by Hershcovich et al. Vijay and Hershcovich (2024) provides an important foundation for cross-cultural NLP by mapping out four main cultural elements: linguistic form, common ground, aboutness and values, to guide researchers through the conceptual complexities of working with cultural diversity. Their approach highlights strategic challenges and high-level pathways for ensuring more fair and representative NLP systems. In contrast, our work expands this foundation by focusing directly on how culture is operationalized in evaluation metrics and benchmark design. While Hershcovich et al. advocate for broad methodological reflection and community engagement, our taxonomy critically dissects actual evaluation practices, exposing how specific methodological choices shape, limit, or obscure cultural complexity. We provide prescriptive strategies for recording ambiguity, contestation, and dissent as core data signals, and for institutionalizing plural ground truths within benchmarks. Therefore, while the Hershcovich et al. framework maps the strategic landscape of the field, our work intervenes more directly at the level of metric design, offering concrete recommendations for embracing the messiness and multiplicity of lived culture in NLP evaluation.
In parallel with our work, Liu et al. Liu et al. (2025b) propose a three-branch taxonomy, i.e., idealistic, linguistic, and social, grounded in anthropology, and used to organize existing NLP resources. Their survey structures elements such as knowledge, values, artifacts, dialects, context, communicative goals, and demographics, emphasizing fine-grained categorization and gaps in coverage.
In contrast, we interrogate how culture is operationalized in the evaluation metrics themselves, making visible methodological practices such as annotation, bias assessment, and consensus building that can flatten cultural complexity. By analyzing how culture is already encoded in evaluation and urging benchmarks to record zones of dispute and ambiguity, our taxonomy moves beyond cataloging toward a design roadmap for resources that reflect multiplicity, contestation, and lived experience.
@@ 3 Culture Benchmark Taxonomy
While technical mitigation strategies have advanced the capture of cultural complexity in NLP, our contribution is to systematically examine how culture itself is conceptualized within leading benchmarks. We begin by offering a taxonomy that clarifies the diverse definitions of culture used across these resources, helping reveal the range of cultural perspectives that inform benchmark design. Building on this taxonomy, we present recommendations in §5 to inform future benchmark development and encourage a more deliberate engagement with cultural nuance.
We first introduce a four-part taxonomy for cultural evaluation in NLP: culture-as-knowledge, culture-as-preference, culture-as-dynamics, and culture-as-bias. The taxonomy is both descriptive and prescriptive. Descriptively, it distills how current benchmarks operationalize culture; prescriptively, it guides future work to combine lenses rather than treat them in isolation, encouraging multidimensional designs that better capture cultural nuance. Appendix B documents the details of the derivation.
@@
Culture consists of facts, traditions, and symbolic references that can be recalled or identified. This lens treats culture as a body of factual or object-level information tied to national or regional identities, for example, benchmarks that assess whether models recognize culturally specific foods, holidays, or practices. This view echoes early trait list approaches in anthropology.
@@
Culture is not treated as static, but as an entity that acts in specific contexts through language, narrative, and interaction. It draws on linguistic and symbolic anthropology to understand culture as enacted, improvised, and contextual, shaped by speaker positionality, audience, and occasion (Goffman, 1959; Hymes, 1974; Bauman and Briggs, 1990). This view understands culture as something that people actively perform in context, rather than an attribute they possess. It focuses on how meaning is shaped through interaction, making it better suited for capturing how models respond to different social situations.
@@
This dimension particularly frames cultures through a series of stereotypes. The goal of this identification is harm reduction and the elimination of bias. It tests how the model outputs reflect and reinforce stereotypes. We introduce this category specifically for NLP evaluations: culture is inferred from patterns of stereotyping or discriminatory output, especially along dimensions of race, gender, or class. Although these benchmarks play an essential role in fairness research, they often rely on binary identity comparisons without interrogating the cultural production of these categories. We aim to introduce an intersectional lens through this category Crenshaw (1991); Alim (2009).

### v2

@@ Abstract
Cultural evaluation of large language models has become increasingly important, yet current benchmarks often reduce culture to static facts or homogeneous values. This view conflicts with anthropological accounts that emphasize culture as dynamic, historically situated, and enacted in practice. In this position paper, we qualitatively examine 20 handpicked cultural benchmarks to show the breadth of how culture is framed within NLP. First, we introduce a four-part framework that categorizes how benchmarks frame culture, such as knowledge, preference, performance, or bias. Using this lens and identify six recurring methodological issues: including treating countries as cultures, overlooking within-culture diversity, and relying on oversimplified survey formats. Drawing on established anthropological methods, we propose concrete improvements: incorporating real-world narratives and scenarios, involving cultural communities in design and validation, and evaluating models in context rather than isolation. Our aim is to guide the creation of cultural benchmarks that move beyond simple recall tasks and more faithfully reflect how models respond to complex cultural situations.
@@
Mai AlKhamissi1thanks:  Equal contribution.   Yunze Xiao [^1] footnotemark: 1   Badr AlKhamissi2   Mona Diab1 1Carnegie Mellon University  2EPFL {malkhami,yunzex,mdiab}@andrew.cmu.edubadr.alkhamissi@epfl.ch
@@ 1 Introduction
Figure 1: Cultural Framing in NLP. Our taxonomy of how culture is framed in NLP evaluation. Each quadrant represents a distinct theoretical lens on culture: defining what it entails, illustrating how it is expressed, and providing two representative benchmarks for each framing.
Large language models (LLMs) are now being implemented in translation systems Zhu et al. (2024), educational tools Sonkar et al. (2024), search engines Ziems et al. (2023), and generative platforms that engage diverse public opinions across cultural, linguistic, and political contexts Ziems et al. (2023). As these models become more central to knowledge production and everyday decision-making, it is crucial to assess their sensitivity to cultural nuanceXiao et al. (2025b). In response, NLP researchers are constructing benchmarks that aim to measure cultural competence: the ability of models to respond appropriately to regionally specific norms, moral frameworks, idioms and socio-political identities.
@@
However, in this growing body of work, the concept of culture is often treated as a background variable rather than as a central analytical concern. Benchmark designers rarely engage with anthropology, a discipline that has developed rich tools to understand how culture is lived, contested, and transmitted. Anthropologists study culture not as a fixed set of values or national traits but as a dynamic and situated process shaped by history, power, and everyday meaning-making Geertz (1973).
@@
To bridge the gap between NLP and anthropology, we analyze 20 cultural benchmarks used in NLP. We argue that while these benchmarks address important questions, such as moral diversity, stereotype reproduction, and regional variation—they often rely on reductive or decontextualized models of culture. Consequently, our contribution is as follows: (1) We examine how culture is conceptualized within benchmarks, developing a taxonomy of cultural framings used by NLP researchers; (2) We show how six common design choices oversimplify culture, reducing it to surface traits or stereotypes, and restricting which forms of cultural knowledge are recognized or ignored; (3) We offer recommendations for building benchmarks that reflect culture as a lived experience.
2 Methods
The choice to analyze 20 benchmarks reflects a methodological commitment to qualitative, humanistic traditions, as opposed to conducting a large-scale survey over many more benchmarks. This (our) approach foregrounds epistemologies, assumptions, and cultural politics often obscured in large-scale corpus analyses. Rather than prioritizing breadth, it emphasizes interpretive depth, reflexivity, and contextual understanding, aligning with scholarship that highlights the limits of scalable methods for capturing nuance, bias, and power dynamics in benchmarks. Our goal is to provide conceptual clarity rather than statistical coverage. Scholars such as Blodgett et al. and voices in ACM interactions (Liao et al., 2024; Widder and Kneese, 2025) stress that meaningful cultural analysis demands slow, interdisciplinary, and critically reflexive engagement, a position this study adopts through its carefully curated sample.
Methodologically, many of these datasets are constructed with expert or native annotator input, guaranteeing high-quality, culturally nuanced data that reflect real-world complexities. This set is chosen because it offers a more holistic and representative coverage of cultural dimensions, while addressing current research gaps in cross-cultural NLP, and is aligned with state-of-the-art evaluation needs. They incorporate rigorous annotation protocols, rendering this curated set a more robust and scientifically justified foundation for comprehensive cultural analysis in NLP.
@@ 3.1 What is Culture
Anthropologists have historically studied culture through immersive qualitative methods such as ethnography, long-term fieldwork, and participant observation. Rather than reducing culture to a set of traits or norms, anthropologists emphasize its fluidity, contestation, and entanglement with everyday practices. Thus, the concept of culture has evolved significantly in the social sciences, shifting from fixed categories to fluid and dynamic interpretations. Early anthropological work, notably Boas’ cultural relativism (Boas, 1911), rejected universal hierarchies and emphasized understanding each culture on its own terms. Boas argued against simplistic "trait-list" models, proposing instead that culture should be understood historically and contextually.
Subsequent interpretive approaches, epitomized by Geertz (1973), further reframed culture as "webs of significance" woven by human beings. Geertz emphasized that cultural analysis requires a "thick description," the detailed interpretation of meanings embedded in everyday actions, rituals, and symbols. This perspective shifted the focus from cultural traits toward the interpretive labor necessary to understand situated symbolic practices. The interpretative approach gave way to the "writing culture" turn, where culture is seen as an object that exists outside of the process of writing it. It is not an object to be discovered "out there," but it is constructed through the process of writing about culture (Clifford and Marcus, 1986). This approach merges researcher with subject, showing that there can be no objective understanding of culture outside the eyes and pen of the researcher.
Building on these foundations, Goffman (1959) introduced the notion of culture as performative and interactional, suggesting that cultural meanings are continuously enacted and negotiated through everyday interactions and social performances. Rather than viewing culture as something individuals inherently possess, this conception emphasizes that culture emerges dynamically in social contexts, shaped by the roles people assume and the interactions they engage in. We use these frameworks to create the taxonomy of culture.
@@ 3.2 Culture Benchmarks in NLP
NLP research has developed benchmarks to evaluate the handling of cultural knowledge using language models, but these works often define culture inconsistently and simplistically compared to anthropological scholarship. Although NLP researchers increasingly recognize that linguistic and social norms vary across cultures and are underrepresented in current resources, significant gaps remain between computational approaches and social science perspectives on culture. The purpose of this work is to foster interdisciplinary dialogue between these fields.
Recent work on cultural and bias evaluation has produced diverse datasets investigating the concept of culture. CultureAtlas Fung et al. (2024) broadens multilingual evaluation, integrating everyday cultural knowledge across diverse global contexts, while SeaExam Liu et al. (2025a) emphasizes reasoning and consistency specifically in Southeast Asian scenarios. Furthermore, CDEval Wang et al. (2024b) systematically maps six cultural dimensions to seven domains, while LLM-GLOBE Karinshak et al. (2024) adapts the well-established GLOBE cultural values framework to evaluate open-ended LLM outputs.
Datasets focused on normative and moral reasoning continue to expand, exemplified by Social Chemistry 101 Forbes et al. (2020), which distills common-sense social norms from real-world interactions; and MoralBench Ji et al. (2024) provides explicit ethical dilemmas for detailed evaluation. Moreover, NormAd Rao et al. (2025) systematically assesses how models judge the acceptability of everyday etiquette scenarios in 75 countries.
Parallel efforts in bias evaluation are significantly diversified: BBQ Parrish et al. (2022) identifies stereotypical biases in question-answering tasks; CrowS-Pairs Nangia et al. (2020) targets implicit biases in masked language models; Social Bias Frames Sap et al. (2020) annotates offensive implications in language model outputs; and BOLD Dhamala et al. (2021) evaluates bias in open-ended biographical generation. The VL-BiasBench Wang et al. (2024a) and multilingual M5 Schneider and Sitaram (2024a) suites extend bias assessment to vision language models, while M-Bias Raza et al. (2024) offers data-driven interventions to reduce toxicity without sacrificing contextual coherence.
Finally, novel datasets are increasingly expanding into affective and application-specific areas: CuLEmo Belay et al. (2025) benchmarks emotion recognition across cultures, CASA Qiu et al. (2025) evaluates social and cultural awareness of LLM-powered web agents, and X-FACTR Jiang et al. (2020) assesses multilingual factual knowledge through a cloze-style fill-in-the-blanks question in 23 typologically diverse languages.
@@ 3.3 Critiques within NLP
Recent work in NLP has exposed fundamental weaknesses in the conceptualization of culture in cultural benchmarks. In the literature, there is a consensus that current evaluation frameworks reduce culture to survey-derived constructs that privilege western-centric assumptions while failing to capture dynamic, situated, and contested cultural meanings Liu (2024a). Hershcovich et al. (2022) identifies unequal cultural representation advocate for targeted sampling and Distributionally Robust Optimization to prevent cultural generalization from being disproportionately shaped by dominant groups. Zhou et al. (2025) extends these concerns by critiquing the over-reliance on nationality-based categorizations and rigid survey methods that flatten dynamic and intersectional cultural identities, proposing more context-sensitive and stereotype-aware datasets to reflect lived cultural realities.
Several influential papers have further articulated these shortcomings. For example, Chiu et al. (2024) highlight how existing benchmarks are often static, relying heavily on sources such as Wikipedia, and are constrained by narrow survey formats that limit the representation of cultural diversity. These practices risk amplifying urban and western-centric narratives while marginalizing underrepresented perspectives. To counteract this, they introduce CulturalBench, a benchmark designed through human-authored and verified questions that span multiple regions and cultural dimensions, with independent annotator validation to improve robustness and inclusivity. In the same vein, Culture is Not Trivia (Zhou et al., 2025) critiques the over-reliance on nationality and rigid survey methods, arguing that such approaches flatten dynamic and intersectional cultural identities. They propose more context-sensitive and stereotype-aware datasets to better reflect lived cultural realities and support inclusive benchmark design.
@@
Researchers have attempted to address these limitations through technical interventions, most notably cultural prompting and semantic augmentation. For example, AlKhamissi et al. (2024) proposed anthropological prompting to make LLMs better reflect the opinions of digitally underrepresented groups. Similarly, Tao et al. (2024) shows that explicitly instructing LLMs to adopt specific cultural perspectives can significantly reduce Western bias in most countries and territories, although this approach sometimes introduces new biases and rests on the problematic assumption that culture can be cleaned up by demographic targeting. Although these innovations show promise in controlled settings, they fundamentally maintain the flawed premise that culture can be distilled through survey data or prompt engineering, an assumption increasingly questioned within the research community.
3.4 Existing Frameworks for Cultural Evaluation
@@
The framework presented by Vijay and Hershcovich (2024) provides an important foundation for cross-cultural NLP by mapping out four main cultural elements: linguistic form, common ground, aboutness, and values, to guide researchers through the conceptual complexities of working with cultural diversity. Their approach highlights strategic challenges and high-level pathways for ensuring more fair and representative NLP systems. In contrast, our work expands this foundation by focusing directly on how culture is operationalized in evaluation metrics and benchmark design.
Concurrent with our work, Adilazuarda et al. (2024) Adilazuarda et al. offer a high-level taxonomy that distinguishes between demographic and semantic proxies of culture, revealing the field’s reliance on operational shortcuts rather than explicit theorizing cultural processes. Complementarily, Liu et al. (2023) Liu (2024b) propose a fine-grained taxonomy that organizes cultural elements into ideational, linguistic, and social dimensions, grounded in anthropological definitions and emphasizing the interplay between language and social practice.
Havaldar et al. (2025) and Pawar (2025) push cultural evaluation beyond static text benchmarks. Havaldar et al. introduce a sociocultural framework for the evaluation of conversational LLM, focusing on linguistic style as a key site where cultural differences manifest, and proposing new metrics such as conversational structure, stylistic sensitivity, and subjective correctness to assess how models adapt across cultural contexts. Pawar (2025) similarly broadens the field by offering a cross-modal taxonomy, connecting textual, visual, and multimodal forms of cultural understanding through dimensions of representation, knowledge, and social sensibility
Together, these frameworks have provided important structure for an emerging interdisciplinary space, but they also highlight the limits of static categorization. Building on this work, the taxonomy advanced here engages both perspectives: acknowledging the functional distinctions emphasized by Adilazuarda et al. (2024) and the theoretical breadth introduced by Liu et al. (2023). We also draw on more recent developments such as Havaldar et al. (2025), who extend cultural analysis into conversational and stylistic metrics by designing a new benchmark, and Pawar (2025), which integrates multimodal dimensions of cultural understanding. Our framework differs from these by centering the conceptual modeling of culture itself rather than its downstream manifestations in text, interaction, or modality. In our conception, culture is not treated as a set of characteristics, but as a shifting interpretive process situated in discourse, interaction, and power, enabling a richer analysis of how LLMs reproduce and transform cultural meaning across contexts.
@@ 4 Culture Benchmark Taxonomy
While technical mitigation strategies have advanced the capture of cultural complexity in NLP, our contribution is to systematically examine how culture itself is conceptualized within leading benchmarks. We begin by offering a taxonomy that clarifies the diverse definitions of culture used across these resources, helping reveal the range of cultural perspectives that inform benchmark design. Building on this taxonomy, we present recommendations in §6 to inform future benchmark development and encourage a more deliberate engagement with cultural nuance. We first introduce a four-part taxonomy for cultural evaluation in NLP: culture-as-knowledge, culture-as-preference, culture-as-dynamics, and culture-as-bias. The taxonomy is both descriptive and prescriptive. Descriptively, it distills how current benchmarks operationalize culture; prescriptively, it guides future work to combine lenses rather than treat them in isolation, encouraging multidimensional designs that better capture cultural nuance.
@@
Culture consists of facts, traditions, and symbolic references that can be recalled or identified. This lens treats culture as a body of factual or object-level information tied to national or regional identities, for example, benchmarks that assess whether models recognize culturally specific foods, holidays, or practices. This view echoes early trait list approaches in anthropology (Boas, 1911).
@@
Culture is not treated as static, but as an entity that acts in specific contexts through language, narrative, and interaction. It draws on linguistic and symbolic anthropology to understand culture as enacted, improvised, and contextual, shaped by speaker positionality, audience, and occasion (Goffman, 1959; Hymes, 1974; Bauman and Briggs, 1990). This view understands culture as something that people actively perform in context, rather than an attribute they possess. It focuses on how meaning is shaped through interaction, making it better suited for capturing how models respond to different social situations. Benchmarks such as CoCo-LM, BLEND, Moral Stories, and CuLEmo exemplify this view by modeling how culture is performed in dialogue, story, or emotion expression.
@@
This dimension particularly frames cultures through a series of stereotypes. The goal of this identification is harm reduction and the elimination of bias. It tests how the model outputs reflect and reinforce stereotypes. We introduce this category specifically for NLP evaluations: culture is inferred from patterns of stereotyping or discriminatory output, especially along dimensions of race, gender, or class. Although these benchmarks play an essential role in fairness research, they often rely on binary identity comparisons without interrogating the cultural production of these categories. We aim to introduce an intersectional lens through this category Crenshaw (1991); Alim (2009). BBQ, Social Bias Frames, BOLD, M-Bias, and VL-Bias Bench all operate within this paradigm, focusing on identity-based harm and fairness violations.
4.1 Contribution
While Hershcovich et al. (2022) advocate for broad methodological reflection and community engagement, our taxonomy critically dissects actual evaluation practices, exposing how specific methodological choices shape, limit, or obscure cultural complexity. We provide prescriptive strategies for recording ambiguity, contestation, and dissent as core data signals, and for institutionalizing plural ground truths within benchmarks. Therefore, while their framework maps the strategic landscape of the field, our work intervenes more directly at the level of metrics, offering concrete recommendations for embracing the messiness and multiplicity of lived culture in NLP evaluation.
In contrast, we interrogate how culture is framed in the evaluation metrics themselves, making visible methodological practices such as annotation, bias assessment, and consensus building that can flatten cultural complexity. By analyzing how culture is already encoded in evaluation and urging benchmarks to record zones of dispute and ambiguity, our taxonomy moves beyond cataloging toward a design roadmap for resources that reflect multiplicity, contestation, and lived experience. .